OpenStudy (anonymous):

Find the distance between each paor of points. Round your answer to the nearest tenth. (-2, -3) and (-4, 4) A) 6.08 B) 4.12 C) 9.22 D) 7.28

OpenStudy (anonymous):

Use the distance formula=root((x2-x1)^2-(y2-y1)^2)=root(4+49)=root(53)=7.28

OpenStudy (aravindg):

do u knw distance formula?

OpenStudy (cwrw238):

yup - this is really Pythagoras theorem

OpenStudy (anonymous):

Not a Pythagoras theorem.Its the distance formula

OpenStudy (anonymous):

You was wrong, its A which is 5.39 units

OpenStudy (anonymous):

What is the value inside the roots u got?

OpenStudy (anonymous):

sorry,in the formula Inside the root its + not a - in between x and y

OpenStudy (cwrw238):

distance = sqrt(4-(-3)^2 + (-4-(-2))^2) = sqrt(7^2 + (-2)^2) = sqrt 53

OpenStudy (cwrw238):

sqrt53 = 7.28

OpenStudy (cwrw238):

if you drew the line on a graph you will see that its the application of pythagoras theorem

OpenStudy (anonymous):

Yeah,agreed

OpenStudy (anonymous):

@Cwrw238 your wrong, it's not 7.28.. It's A